Solve the system of linear equations by adding or subtracting.

−2x + y = −7
2x − 6y = −8
Enter the solution as an ordered pair.

The solution is?

Answer by rumessa(13)

by adding both equation
(-2x +y = -7) + ( 2x - 6y = -8)
-5y = -15
y= 3
put the value of y in equation 1
-2x + 3 = -7
-2x = -10
x= 5
(x,y) = (5,3)